# Existence proofs of traveling wave solutions on an infinite strip for the suspension bridge equation and proof of orbital stability

### v1

@@ 1 Introduction
The objective of this paper is to prove the existence of traveling wave solutions of the two-dimensional suspension bridge equation on an infinite strip $\Omega\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}{\mathbb{R}}\times(-d_{2},d_{2})$ for $d_{2}>0$. We let $v=v(t,X_{1},X_{2}):{\mathbb{R}}^{+}\times\Omega\to{\mathbb{R}}$ represent the deflection in the downward direction of the surface of the bridge and we let $\Delta^{2}=(\partial_{X_{1}}^{2}+\partial_{X_{2}}^{2})^{2}$. One of the systems used for modeling suspension bridges is (e.g. [8], [18])
@@
We only consider values of $c$ in $(0,\sqrt{2})$, because if $c$ tends to zero, the amplitude of the waves grows closer to infinity, and for $c=0$, nontrivial solutions do not longer exist [23]. Contrarily, if $c$ approaches $\sqrt{2}$, the solutions become oscillatory as the amplitudes of the waves tend to go to zero [15].
Our goal is to establish the existence of a solution $\tilde{u}$ to (2) on $\Omega$. Numerical simulations of solutions to this equation, obtained via the Mountain Pass Algorithm, are discussed in [15]. A rigorous existence proof for traveling wave solutions to (2) on finite domains, periodic in one direction and satisfying Neumann boundary conditions in the other, was later given in [32]. We combine the techniques and results outlined there with procedures developed in [6] and [7] to obtain an existence proof of a solution that decays to zero as $|x_{1}|\rightarrow\infty$.
@@
Ultimately our goal is to construct an approximate solution $\bar{u}:\Omega\to{\mathbb{R}}$ to (2) and then prove the existence of a true solution to (2) in a vicinity of $\bar{u}$. An illustration of such an approximate solution $\bar{u}$, that we rigorously verify in Theorem 6.1, is depicted in Figure 1. The precise construction of $\bar{u}$ is explained in Section 3.1. The verification of such an approximate solution is done using a computer-assisted proof (CAP), a method that has gained significant interest in the rigorous verification of numerical simulations for dynamical systems over the past few decades. Early implementations of CAPs to the study of elliptic PDEs are seen in [19] and [24], stemming from 1988 and 1992, respectively. Later, advancements in the applications of CAPs to (elliptic) PDEs were made (e.g. [9, 10, 21, 27, 30, 34]). Comprehensive discussions on the role of CAPs in the study of PDEs are outlined in [14] and [20].
In particular, the setting in which we work allows us to utilize a Newton-Kantorovich-type theorem. These type of theorems are widely used in the construction of CAPs for PDEs and various other contexts (see for instance [1, 16, 29]). Specifically, we use a variant of this theorem called the radii polynomial theorem [10]. To compensate for the fact that we are working on an unbounded domain, an additional bound that needs to be computed is introduced, on top of the other bounds mentioned in the radii polynomial theorem. To guarantee the reliability of the computational components of the proof, we employ Julia interval arithmetic using the packages [17] and [26].
The solutions obtained in the CAPs will also be analyzed for stability. This part of the work is based on the procedure outlined in [18], where the authors investigate orbital stability of one-dimensional traveling wave solutions to the suspension bridge equation. Their analysis employs energy methods and spectral analysis to establish the stability results, supported by computer-assisted techniques. In two dimensions, to our knowledge, only numerical results concerning stability are available so far, which can be found in [15]. To rigorously handle the spectral properties of the higher-dimensional equation, we follow the framework proposed in [4]. Using this approach, we are able to conclude that the solution corresponding to the approximation in Figure 1 is orbitally stable.
@@ 2.1 Zero-finding problem for functions
$L^{2}_{e}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{u(x_{1},x_{2})=u_{0}(x_{1})+2\sum_{n_{2}\in\mathbb{N}}u_{n_{2}}(x_{1})\cos\left(2\pi\tilde{n}_{2}x_{2}\right),~u_{n_{2}}(x)=u_{n_{2}}(-x)\text{ and }\|u\|_{2}<\infty\right\}$
@@
Now, it remains to prove that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is well defined in order to set-up (2) on $\mathcal{H}$. For that purpose, we use the following result:
@@
We conclude the proof using (57) in Lemma 7.1. ∎
@@
In practice, for the computation of $\kappa_{2}$, we use Riemann summations to bound the summation up to finitely many $n_{2}$. In fact, this can be achieved using rigorous numerics (cf. LA: Reference code). To estimate the tail of the summation, we use (58) in Lemma 7.1.
Using the above lemma, we have that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is a smooth operator. Accordingly, we investigate the solution of the following zero-finding problem
@@ 2.2 Fourier series representation
As described in [6], existence proofs of localized solutions can be achieved thanks to approximate objects on Fourier coefficients. Following such an approach, we introduce notation corresponding to objects on Fourier coefficients, related to the ones introduced in the previous section. Let $d_{1}>0$ and define $\Omega_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(-d_{1},d_{1})\times(-d_{2},d_{2})$. $\Omega_{0}\subset\Omega$ is the bounded domain on which our approximate solution is constructed. We let $\mathbb{N}_{0}^{2}=(\mathbb{N}\cup\{0\})^{2}$ and we define
@@
$\left((\gamma(u)\right)_{n}$ $=\frac{1}{|\Omega_{0}|}\int_{\Omega_{0}}u(x_{1},x_{2})e^{-2\pi i\tilde{n}\cdot x}dx\quad\text{ for }n\in\mathbb{N}_{0}^{2},$
@@
for all bounded linear operators $\mathbb{H}\in\mathcal{B}(L^{2}_{e})$ and $H\in\mathcal{B}(\ell^{2})$. The map $\Gamma^{\dagger}$ will be useful when constructing operators on $\mathcal{B}(L^{2}_{e})$ thanks to operators on Fourier coefficients (that is, operators on $\mathcal{B}(\ell^{2})$). Furthermore, we introduce the norm $\|\cdot\|_{2,\mathcal{H}}$ to denote the operator norm for bounded linear operators on $L^{2}_{e}\rightarrow\mathcal{H}$.
@@ 3 Constructive proof of existence
In this section, we present our approach for proving constructively the existence of solutions to (2). Our approach is based on a Newton-Kantorovich theorem. More specifically, we construct a fixed point operator for which fixed points correspond to zeros of $\mathbb{F}$ (8).
First, we construct an approximate solution $\bar{u}\in\mathcal{H}$ thanks to its Fourier coefficients representation. The technical construction of $\bar{u}$ is described in Section 3.1. Given $\bar{u}$, we then construct $\mathbb{A}:L^{2}_{e}\to\mathcal{H}$ as an approximate inverse for the Fréchet derivative $D\mathbb{F}(\bar{u})$. Then, our goal is to determine a radius $r>0$ such that $\mathbb{T}$, given as
@@
is well defined on $\overline{B_{r}(\bar{u})}\to\overline{B_{r}(\bar{u})}$ and contracting (where $B_{r}(\bar{u})$ is the open ball of $\mathcal{H}$ centered at $\bar{u}$ and of radius $r$). The Banach fixed point theorem then provides the existence of a unique zero $\tilde{u}\in\mathcal{H}$ in $\overline{B_{r}(\bar{u})}.$
@@ 3.1 Construction of an approximate solution
Now, note that by construction $\bar{u}\in L^{2}_{e}$ but we do not necessarily have $\bar{u}\in\mathcal{H}.$ Indeed, $\bar{u}$ might not be smooth at $x_{1}=\pm d_{1}$. By periodicity of $\bar{u}$ on $\Omega_{0}$}, this is equivalent for asking smoothness only at $d_{1}$. Now, in order to obtain the required regularity for $\mathcal{H}$, we need
@@
Note that since $\bar{u}$ has a cosine-cosine representation, we readily have $\partial_{x_{1}}^{k}\bar{u}(d_{1},x_{2})=0$ for all $x_{2}\in(-d_{2},d_{2})$ and all $k\in\{1,3\}$. Consequently, it remains to ensure that
@@
Note that since $\bar{u}$ is represented thanks to $\bar{U}$ on $\Omega_{0}$}, (19) can be translated as equations on $\bar{U}$ :
@@
where we have used that $\bar{U}=\pi^{N}\bar{U}$. Note that this is equivalent to having $\bar{U}\in Ker(\mathcal{T})$, where
@@
Consequently, $\bar{U}$ needs to be projected into the kernel of the matrix $\mathcal{T}$. Such an operation has been described in [6, 7] and is implemented in [31].
Overall, we ensure that $\bar{U}\in Ker(\mathcal{T})$ and obtain that
@@ 3.2 Computing the nonlinearity
$e^{\bar{U}}\in\begin{cases}{(e^{\bar{U}})}^{\mathrm{FFT}}_{n}+{C}[-\varepsilon_{n},\varepsilon_{n}]&\text{ if }0\leq n_{1}\leq N^{0}_{1},0\leq n_{2}\leq N^{0}_{2},\\ [-\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}},\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}}]&\text{ otherwise}\end{cases}$ (21)
@@ 4 Computation of the bounds
In this section, we justify our choices for the approximate objects constructed above by providing explicit formulas for the bounds of Theorem 3.1.
@@
Recall that $N^{0}$ corresponds to the truncation dimension for sequences and $N$ for operators. Note that the nonlinear part of $F(\bar{U})$ in $\|B^{N}F(\bar{U})\|_{2}^{2}$ can be calculated using (21).
@@
$\|(\pi^{N^{0}}-\pi^{N})(L\bar{U}+G(\bar{U}))\|_{2}^{2}\in\sum_{\begin{subarray}{c}n\in I_{N^{0}}\\ n\notin I_{N}\end{subarray}}\alpha_{n}\left(l(\tilde{n})(\bar{U})_{n}+(e^{\bar{U}})^{\mathrm{FFT}}+C[-\varepsilon_{n},\varepsilon_{n}]\right)^{2}.$
@@
First, we recall the computation of the bound $Z_{1}$, which is given in Lemma 3.9 from [7].
@@
$\mathcal{Z}_{u,1}$ $\leq 2d_{1}(\bar{V}^{N},\bar{V}^{N}*E)_{2},$
$\mathcal{Z}_{u,2}$ $\leq 2d_{1}(\bar{V}^{N},V^{N}*E)_{2}+C(d_{1})e^{-2ad_{1}}(\bar{V}^{N},\bar{V}^{N}*E)_{2},$
@@
for all $n_{2}\in\mathbb{N}_{0}$, where we have used (28). Using that $v=\bar{v}^{N}u$, we have
@@
Now, using the Cauchy-Schwarz inequality, we get
@@
since $\|u\|_{2}=1.$ Now, using Fubini’s theorem, we have
@@
Using the proof of Theorem 3.9 from [6] again, we get
@@
for all $n_{1}\in\mathbb{N}_{0}$ and all $n_{2}\in\mathbb{N}_{0}$. Using Parseval’s identity, this implies that
@@
This concludes the proof for $\mathcal{Z}_{u,1}$.
For the computation of $\mathcal{Z}_{u,2}$ we use the formulas provided in [6]. Indeed, using the proof of Theorem 3.9, we can define $g_{n}$ as
@@
for all $n\in\mathbb{Z}^{2}$, where $v_{n_{2}}=v_{-n_{2}}$ if $n_{2}<0$. Using the proof of Theorem 3.9 in [6], we have that
@@
Using the proof of Lemma 6.5 in [6], we obtain that

### v3

# Existence and orbital stability proofs of traveling wave solutions on an infinite strip for the suspension bridge equation

@@ 1 Introduction
The objective of this paper is to prove the existence of traveling wave solutions of the two-dimensional suspension bridge equation on an infinite strip $\Omega\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}{\mathbb{R}}\times(-d_{2},d_{2})$ for $d_{2}>0$ and to prove orbital (in)stability. We let $v=v(t,X_{1},X_{2}):{\mathbb{R}}^{+}\times\Omega\to{\mathbb{R}}$ represent the deflection in the downward direction of the surface of the bridge and we let $\Delta^{2}=(\partial_{X_{1}}^{2}+\partial_{X_{2}}^{2})^{2}$. One of the systems used for modeling suspension bridges is (e.g. [8], [18])
@@
We only consider values of $c$ in $(0,\sqrt{2})$, because if $c$ tends to zero, the amplitude of the waves grows to infinity, and for $c=0$, nontrivial solutions do not longer exist [23]. Contrarily, if $c$ approaches $\sqrt{2}$, the solutions become oscillatory as the amplitudes of the waves tend to go to zero [16].
We want to establish the existence of a solution $\tilde{u}$ to (2) on $\Omega$. Numerical simulations of solutions to this equation, obtained via the Mountain Pass Algorithm, are discussed in [16]. A rigorous existence proof for traveling wave solutions to (2) on finite domains, periodic in one direction and satisfying Neumann boundary conditions in the other, was later given in [32]. We combine the techniques and results outlined there with procedures developed in [4] and [5] to obtain an existence proof of a solution that decays to zero as $|x_{1}|\rightarrow\infty$.
@@
In this paper, we construct an approximate solution $\bar{u}:\Omega\to{\mathbb{R}}$ to (2) and then prove the existence of a true solution to (2) in a vicinity of $\bar{u}$. An illustration of such an approximate solution $\bar{u}$, that we rigorously verify in Theorem 6.1, is depicted in Figure 1. The precise construction of $\bar{u}$ is explained in Section 3.1. The verification of such an approximate solution is done using a computer-assisted proof (CAP), a method that has gained significant interest in the rigorous verification of numerical simulations for dynamical systems over the past few decades. Early implementations of CAPs to the study of elliptic PDEs are seen in [20] and [24], stemming from 1988 and 1992, respectively. Later, advancements in the applications of CAPs to (elliptic) PDEs were made (e.g. [9, 10, 21, 27, 30, 34]). Comprehensive discussions on the role of CAPs in the study of PDEs are outlined in [14] and [19].
In particular, the setting in which we work allows us to utilize a Newton-Kantorovich-type theorem. These type of theorems are widely used in the construction of CAPs for PDEs and various other contexts (see for instance [1, 17, 29]). Specifically, we use a variant of this theorem called the radii polynomial theorem [10]. To prove existence on a bounded domain using this theorem, we need to compute certain bounds. To compensate for the fact that we are working on an unbounded domain, an additional bound is introduced and computed in this paper. To guarantee the reliability of the computational components of the proof, we employ Julia interval arithmetic using the packages [15] and [26].
The solutions obtained in our CAPs will also be analyzed for stability. This part of the work is based on the procedure outlined in [18], where the authors investigate orbital stability of one-dimensional traveling wave solutions to the suspension bridge equation. Their analysis employs energy methods and spectral analysis to establish the stability results, supported by computer-assisted techniques. In two dimensions, to our knowledge, only numerical results concerning stability are available so far, which can be found in [16]. To rigorously handle the spectral properties of the higher-dimensional equation, we follow the framework proposed in [7]. Using this approach, we are able to conclude that the solution corresponding to the approximation in Figure 1 is orbitally stable.
@@ 2.1 Zero-finding problem for functions
$L^{2}_{e}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\left\{u(x_{1},x_{2})=u_{0}(x_{1})+2\sum_{n_{2}\in\mathbb{N}}u_{n_{2}}(x_{1})\cos\left(2\pi\tilde{n}_{2}x_{2}\right),~u_{n_{2}}(x_{1})=u_{n_{2}}(-x_{1})\text{ and }\|u\|_{2}<\infty\right\}$
@@
Now, it remains to prove that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is well defined in order to formulate (2) on $\mathcal{H}$. For that purpose, we use the following result:
@@
We conclude the proof using (57) from Lemma 8.1. ∎
@@
In practice, for the computation of $\kappa_{2}$, we use Riemann summations to bound the summation up to finitely many $n_{2}$. In fact, this can be achieved using rigorous numerics (cf. [31]). To estimate the tail of the summation, we use (58) in Lemma 8.1.
Using the above lemma, we have that $\mathbb{G}:\mathcal{H}\to L^{2}_{e}$ is a smooth function. Accordingly, we investigate the solution of the following zero-finding problem
@@ 2.2 Fourier series representation
As described in [4], existence proofs of localized solutions can be achieved using approximations of the true solution in Fourier space. Following such an approach, we introduce notation corresponding to Fourier series, related to the ones introduced in the previous section. Let $d_{1}>0$ and let $\Omega_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(-d_{1},d_{1})\times(-d_{2},d_{2})$ be the bounded domain on which our approximate solution is constructed. We let $\mathbb{N}_{0}^{2}=(\mathbb{N}\cup\{0\})^{2}$ and we define
@@
$\left(\gamma(u)\right)_{n}$ $=\frac{1}{|\Omega_{0}|}\int_{\Omega_{0}}u(x_{1},x_{2})e^{-2\pi i\tilde{n}\cdot x}dx\quad\text{ for }n\in\mathbb{N}_{0}^{2},$
@@
for all bounded linear operators $\mathbb{H}\in\mathcal{B}(L^{2}_{e})$ and $H\in\mathcal{B}(\ell^{2})$. The map $\Gamma^{\dagger}$ will be useful when constructing operators on $\mathcal{B}(L^{2}_{e})$ thanks to operators on Fourier coefficients (that is, operators on $\mathcal{B}(\ell^{2})$). Furthermore, we introduce the norm $\|\cdot\|_{2,\mathcal{H}}$ to denote the operator norm for bounded linear operators $L^{2}_{e}\rightarrow\mathcal{H}$.
@@ 3 Constructive proof of existence
In this section, we present our approach for proving constructively the existence of solutions to (2). Our approach is based on a Newton-Kantorovich theorem. More specifically, we construct a fixed point operator for which fixed points correspond to zeros of $\mathbb{F}$.
First, we construct an approximate solution $\bar{u}\in\mathcal{H}$ thanks to its Fourier coefficients representation. The technical construction of $\bar{u}$ is described in Section 3.1. Given $\bar{u}$, we then construct an injective $\mathbb{A}:L^{2}_{e}\to\mathcal{H}$ as an approximate inverse for the Fréchet derivative $D\mathbb{F}(\bar{u})$. Then, our goal is to determine a radius $r>0$ such that $\mathbb{T}$, given as
@@
is well defined on $\overline{B_{r}(\bar{u})}\to\overline{B_{r}(\bar{u})}$ and contracting (where $B_{r}(\bar{u})$ is the open ball of $\mathcal{H}$ centered at $\bar{u}$ and of radius $r$). The Banach fixed point theorem then provides the existence of a unique zero of $\tilde{u}\in\mathcal{H}$ of $\mathbb{F}$ in $\overline{B_{r}(\bar{u})}.$
@@ 3.1 Construction of an approximate solution
Now, note that by construction $\bar{u}\in L^{2}_{e}$, but we do not necessarily have $\bar{u}\in\mathcal{H}.$ Indeed, $\bar{u}$ might not be smooth at $x_{1}=\pm d_{1}$. By periodicity of $\bar{u}$ on $\Omega_{0}$}, smoothness at $x_{1}=\pm d_{1}$ is equivalent to smoothness only at $d_{1}$. Now, in order to obtain the required regularity for $\mathcal{H}$, we need
@@
Since $\bar{u}$ has a cosine-cosine representation, we readily have $\partial_{x_{1}}^{k}\bar{u}(d_{1},x_{2})=0$ for all $x_{2}\in(-d_{2},d_{2})$ and all $k\in\{1,3\}$. Consequently, it remains to ensure that
@@
As $\bar{u}$ is represented by $\bar{U}$ on $\Omega_{0}$}, (19) can be translated to equations on $\bar{U}$ :
@@
where we have used that $\bar{U}=\pi^{N}\bar{U}$. Note that this is equivalent to having $\bar{U}\in\mathrm{Ker}(\mathcal{T})$, where
@@
Consequently, $\bar{U}$ needs to be projected into the kernel of the matrix $\mathcal{T}$ of size $(N^{0}_{1}+1)(N^{0}_{2}+1)$. Such an operation has been described in [4, 5] and is implemented in [31].
Overall, we ensure that $\bar{U}\in\mathrm{Ker}(\mathcal{T})$ and obtain that
@@ 3.2 Computing the nonlinearity
$\left(e^{\bar{U}}\right)_{n}\in\begin{cases}{(e^{\bar{U}})}^{\mathrm{FFT}}_{n}+{C}[-\varepsilon_{n},\varepsilon_{n}]&\text{ if }0\leq n_{1}\leq N^{0}_{1},0\leq n_{2}\leq N^{0}_{2},\\ [-\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}},\frac{{C}}{\nu_{1}^{|n_{1}|}\nu_{2}^{|n_{2}|}}]&\text{ otherwise}\end{cases}$ (21)
@@ 4 Computation of the bounds
In this section, we justify our choices for the approximate objects constructed above by providing explicit formulas for the bounds of Theorem 3.1. Before stating the first bound, we recall that $N^{0}$ corresponds to the truncation dimension for sequences and $N$ for operators.
@@
Note that the nonlinear part of $F(\bar{U})$ in $\|B^{N}F(\bar{U})\|_{2}^{2}$ can be calculated and bounded using (21).
@@
$\|(\pi^{N^{0}}-\pi^{N})(L\bar{U}+G(\bar{U}))\|_{2}^{2}\in\sum_{\begin{subarray}{c}n\in I_{N^{0}}\\ n\notin I_{N}\end{subarray}}\alpha_{n}\left(l\left(\frac{n}{2d}\right)\bar{U}_{n}+(e^{\bar{U}})_{n}^{\mathrm{FFT}}+C[-\varepsilon_{n},\varepsilon_{n}]\right)^{2}.$
@@
We recall the computation of the bound $Z_{1}$, which is given in Lemma 3.9 from [5].
@@
$\mathcal{Z}_{u,1}$ $\leq\bigg(2d_{1}(\bar{V}^{N},\bar{V}^{N}*E)_{2}\bigg)^{\frac{1}{2}},$
$\mathcal{Z}_{u,2}$ $\leq\bigg(2d_{1}(\bar{V}^{N},V^{N}*E)_{2}+C(d_{1})e^{-2ad_{1}}(\bar{V}^{N},\bar{V}^{N}*E)_{2}\bigg)^{\frac{1}{2}},$
@@
for all $n_{2}\in\mathbb{N}_{0}$, where we have used (28). Noting that $v=\bar{v}^{N}u$, we have
@@
Now, applying the Cauchy-Schwarz inequality, we get
@@
since $\|u\|_{2}=1.$ Employing Fubini’s theorem, we have
@@
Recalling the proof of Theorem 3.9 from [4], we get
@@
for all $n_{1}\in\mathbb{N}_{0}$ and all $n_{2}\in\mathbb{N}_{0}$. Applying Parseval’s identity, this implies that
@@
This concludes the inequality for $\mathcal{Z}_{u,1}$.
For the computation of $\mathcal{Z}_{u,2}$ we use the formulas provided in [4]. Indeed, applying the proof of Theorem 3.9 again, we can define $g_{n}$ as
@@
for all $n\in\mathbb{Z}^{2}$, where $v_{n_{2}}=v_{-n_{2}}$ if $n_{2}<0$. Utilizing the proof of Theorem 3.9 in [4], we have that
@@
With the help of the proof of Lemma 6.5 in [4], we obtain that